Holt Enterprises recently paid a dividend, Do, of $2.00. It expects to have nonconstant growth of 25% for 2 years followed byq 4
Answer #1

1)

The terminal, or horizon, date is the date when the growth rate becomes constant. This occurs at the end of time 2.

2)

Year 1 dividend = 2 * (1 + 25%) = 2.5

Year 2 dividend = 2.5 * (1 + 25%) = 3.125

Year 3 dividend = 3.125 * (1 + 8%) = 3.375

Horizon value = D3 / required rate - growth rate

Horizon value = 3.375 / 0.11 - 0.08

Horizon value = 3.375 / 0.03

Horizon value = $112.50

3)

Intrinsic value = 2.5 / (1 + 0.11)1 + 3.125 / (1 + 0.11)2 + 112.5 / (1 + 0.11)2

Intrinsic value = $96.10